# Superluminal constraints from ultra–high–energy neutrino events

### v1

@@ V.1 Constraints from the KM3–230213A event
    with the precise value depending on the assumed source redshift (higher $z$ implies stronger bounds due to cosmological expansion). These constraints provide the current limits on superluminal neutrino propagation consistent with the observed stability of the KM3–230213A event, and include both VPE and NSpl, their thresholds, and the proper cosmological propagation.
V.2 Compatibility with time–of–flight effects
The implications of Lorentz violation in decay kinematics are closely tied to its possible manifestations in time–of–flight effects. A superluminal dispersion relation implies both (i) a finite decay width at high energy, which threatens the stability of UHE neutrinos over cosmological distances, and (ii) an advance in the arrival time with respect to photons or gravitons from the same transient source. These two aspects have so far been probed separately: on the one hand, searches for neutrino–photon temporal correlations in association with gamma–ray bursts and other transients constrain possible time–of–flight advances; while on the other hand stability arguments applied to the recent KM3–230213A event constrain the allowed decay width at $\sim 100$ PeV. As emphasized in Ref. [21], a combined analysis of both effects for the same event would provide a powerful internal consistency test of LIV. Following this reference, the arrival–time advance relative to a photon can be written, as
@@ V.2 Compatibility with time–of–flight effects
$H_{0}^{-1}\simeq 4.4\times 10^{17}\,$s is the Hubble time, and the redshift integral is of order unity for $z\!\sim\!1$, reaching values $\sim 8$ for $n=2$ and $z\!\sim\!3$ [21].
The bounds over the scale of new physics obtained in the previous section from the stability analysis, can be translated to constraints for the time delay:
  • Case $n=0$: For $\delta\!\sim\!10^{-19}$$10^{-23}$, one finds
@@
  • Case $n=2$: For $\Lambda\!\sim\!(3\times 10^{-2}\text{--}5\times 10^{1})E_{\rm Pl}$ and $E_{d}\!=\!100$ PeV, the associated time advance is
@@
    depending on $\Lambda$ and $z$. Since the bound on $\Lambda$ tightens at high $z$, the corresponding $\Delta t_{2}$ decreases, keeping the effect below the sub-second level across the parameter space favored by the stability constraint. In this scenario, both the decay width and the time advance scale as $E^{2}/\Lambda^{2}$. The stability of the 100 PeV KM3–230213A neutrino already constrains $\Lambda$ to values up to tens of times the Planck scale, so future detections at even higher energies will mainly serve to reinforce these bounds rather than to reach new energy scales. As in the $n=0$ case, timing measurements at lower energies provide a crucial cross–check: their consistency with null time delays would confirm that the same LIV parameters governing decay kinematics also control propagation.
A mismatch between stability and timing constraints—namely, a significant time delay incompatible with the decay–based bounds—would signal a departure from the EFT description of LIV. In such a case, the linear scenario ($n=1$), where neutrinos and antineutrinos receive LIV corrections of opposite sign, would become particularly relevant: if the underlying dynamics lies beyond the EFT regime, the symmetry between superluminal and subluminal behavior could be broken, giving rise to effects distinguishable from those of the standard framework.
In summary, the KM3–230213A event already constrains superluminal propagation to a level where any associated time–of–flight effect is expected to lie well below present detection thresholds. Future UHE neutrino observations, especially in combination with precise timing studies of lower–energy events, will provide a stringent internal consistency test of Lorentz invariance, probing whether the underlying mechanism conforms to an effective field theory description or requires a more fundamental extension.
@@ VI Conclusions and future prospects
In this work we have revisited the constraints on Lorentz Invariance Violation (LIV) that can be derived from the detection of ultra–high–energy (UHE) neutrinos, using as reference the KM3–230213A event recently observed by the KM3NeT collaboration. We have presented a unified and self–consistent framework for superluminal neutrino decays, valid for both the energy–independent ($n=0$) and quadratic ($n=2$) dispersion–relation scenarios. The corresponding decay widths and threshold energies have been consistently implemented, together with the proper cosmological propagation. We have shown that the simplified method based on the survival probability provides a robust approximation for setting LIV bounds, since cascade regeneration effects are suppressed by the steep energy dependence of the decay width and the small survival probabilities used in the limit extraction.
@@
By contrast, the source spectrum does play an important role when going beyond the existence of a single detected event and addressing the predicted flux. In particular, it controls the relative weight of regenerated neutrinos produced in superluminal cascades, and it determines the expected cosmogenic contribution at the highest energies. Thus, while the LIV bounds themselves are largely spectrum–independent, any attempt to model the population of UHE neutrinos and its origin (e.g. cosmogenic vs. source–accelerated) will require assumptions about the injection spectrum, spectral index, and high–energy cutoff.
A second important aspect is the possible distinction between neutrinos and antineutrinos at ultra–high energies. The linear case ($n=1$) differs qualitatively from the $n=0$ and $n=2$ scenarios, since for $n=1$ the sign of the LIV parameter alternates between particles and antiparticles: superluminal neutrinos imply subluminal antineutrinos and vice versa. Determining whether the detected UHE flux is dominated by neutrinos, antineutrinos, or a mixture of both will thus be crucial to extend these analyses to the linear LIV case.
@@
The detection of UHE neutrinos opens a powerful avenue to test Lorentz invariance at unprecedented energies. The next years will likely bring multiple detections of UHE neutrino events by both IceCube and KM3NeT, offering a unique opportunity to cross–validate of energy scales, flavor composition, and sky coverage. A combined analysis of their observations could significantly improve the sensitivity to LIV, as well as to the spectral and directional properties of the UHE neutrino population. The present study establishes a consistent foundation for these analyses, demonstrating that the survival–probability approximation is adequate for setting current bounds, while also outlining the conditions under which cascade effects, flavor composition, or spectral uncertainties may become significant in future, more precise observations.
@@ Acknowledgments
This work is supported by the Spanish grants PGC2022-126078NB-C21, funded by MCIN/AEI/10.13039/501100011033 and ‘ERDF A way of making Europe’, grant E21_23R funded by the Aragon Government and the European Union, and the NextGenerationEU Recovery and Resilience Program on ‘Astrofísica y Física de Altas Energías’ CEFCA-CAPA-ITAINNOVA. The authors would like to acknowledge the contribution of the COST Actions CA18108 “Quantum gravity phenomenology in the multi-messenger approach” and CA23130 “Bridging high and low energies in search of quantum gravity”.

### v2

@@ V.1 Constraints from the KM3-230213A event
    with the precise value depending on the assumed source redshift (higher $z$ implies stronger bounds due to cosmological expansion). These constraints provide the current limits on superluminal neutrino propagation consistent with the observed stability of the KM3-230213A event, and include both VPE and NSpl, their thresholds, and the proper cosmological propagation.
V.2 Compatibility with time-of-flight effects
The implications of Lorentz violation in decay kinematics are closely tied to its possible manifestations in time-of-flight effects. A superluminal dispersion relation implies both (i) a finite decay width at high energy, which threatens the stability of UHE neutrinos over cosmological distances, and (ii) an advance in the arrival time with respect to photons or gravitons from the same transient source. These two aspects have so far been probed separately: on the one hand, searches for neutrino-photon temporal correlations in association with gamma-ray bursts and other transients constrain possible time-of-flight advances; while on the other hand stability arguments applied to the recent KM3-230213A event constrain the allowed decay width at $\sim 100$ PeV. As emphasized in ref. [29], a combined analysis of both effects for the same event would provide a powerful internal consistency test of LIV. Following this reference, the arrival-time advance relative to a photon can be written, as
@@ V.2 Compatibility with time-of-flight effects
$H_{0}^{-1}\approx 4.4\times 10^{17}\,$s is the Hubble time, and the redshift integral is of order unity for $z\!\sim\!1$, reaching values $\sim 8$ for $n=2$ and $z\!\sim\!3$ [29].
The bounds over the scale of new physics obtained in the previous section from the stability analysis, can be translated to constraints for the time delay. Taking $E_{d}=100\,$PeV as before, we have:
  • Case $n=0$: For $\delta\!\sim\!10^{-23}$$10^{-19}$, one finds
@@
  • Case $n=2$: For $\Lambda\!\sim\!(3\times 10^{-2}\text{--}5\times 10^{1})E_{\rm Pl}$, the associated time advance is
@@
    depending on $\Lambda$ and $z$. In this scenario, both the decay width and the time advance scale as $E^{2}/\Lambda^{2}$. As a result, the stability of the $\approx 100\,\mathrm{PeV}$ KM3–230213A neutrino already enforces that any superluminal time advance remains very small. Although increasing the neutrino energy would enhance $\Delta t_{2}$ for a fixed $\Lambda$, the stability requirement simultaneously pushes the allowed values of $\Lambda$ upward, so that the maximal compatible time advance remains at the sub-second level. This implies that, as in the $n=0$ case, timing measurements at lower energies already provide a crucial cross-check: their consistency with null time delays would confirm that the same LIV parameters governing decay kinematics also control propagation.
A mismatch between stability and timing constraints—namely, a significant time delay incompatible with the decay-based bounds—would signal a departure from the EFT description of LIV. In such a case, the linear scenario ($n=1$), where neutrinos and antineutrinos receive LIV corrections of opposite sign, would become particularly relevant: if the underlying dynamics lies beyond the EFT regime, the symmetry between superluminal and subluminal behavior could be broken, giving rise to effects distinguishable from those of the standard framework.
In summary, the KM3-230213A event already constrains superluminal propagation to a level where any associated time-of-flight effect is expected to lie well below present detection thresholds. Future UHE neutrino observations, especially in combination with precise timing studies of lower-energy events, will provide a stringent internal consistency test of Lorentz invariance, probing whether the underlying mechanism conforms to an effective field theory description or requires a more fundamental extension.
@@ VI Conclusions and future prospects
In this work we have revisited the constraints on Lorentz Invariance Violation (LIV) that can be derived from the detection of ultra-high-energy (UHE) neutrinos, using as reference the KM3-230213A event recently observed by the KM3NeT collaboration. We have presented a unified and self-consistent framework for superluminal neutrino decays, valid for both the energy-independent ($n=0$) and quadratic ($n=2$) dispersion-relation scenarios. The corresponding decay widths and threshold energies have been consistently implemented, together with the proper cosmological propagation. We have shown that the simplified method based on the survival probability provides a robust approximation for setting LIV bounds, since cascade regeneration effects are suppressed by the steep energy dependence of the decay width and the small survival probabilities used in the limit extraction.
@@
By contrast, the source spectrum does play an important role when going beyond the existence of a single detected event and addressing the predicted flux. In particular, it controls the relative weight of regenerated neutrinos produced in superluminal cascades, and it determines the expected cosmogenic contribution at the highest energies. Thus, while the LIV bounds themselves are largely spectrum-independent, any attempt to model the population of UHE neutrinos and its origin (e.g. cosmogenic vs. source-accelerated) will require assumptions about the injection spectrum, spectral index, and high-energy cutoff.
A second important aspect is the possible distinction between neutrinos and antineutrinos at ultra-high energies. The linear case ($n=1$) differs qualitatively from the $n=0$ and $n=2$ scenarios, since for $n=1$ the sign of the LIV parameter alternates between particles and antiparticles: superluminal neutrinos imply subluminal antineutrinos and vice versa. Determining whether the detected UHE flux is dominated by neutrinos, antineutrinos, or a mixture of both will thus be crucial to extend these analyses to the linear LIV case.
@@
The detection of UHE neutrinos opens a powerful avenue to test Lorentz invariance at unprecedented energies. The next years will likely bring multiple detections of UHE neutrino events by both IceCube and KM3NeT, offering a unique opportunity to cross-validate of energy scales, flavor composition, and sky coverage. A combined analysis of their observations could significantly improve the sensitivity to LIV, as well as to the spectral and directional properties of the UHE neutrino population. The present study establishes a consistent foundation for these analyses, demonstrating that the survival-probability approximation is adequate for setting current bounds, while also outlining the conditions under which cascade effects, flavor composition, or spectral uncertainties may become significant in future, more precise observations.
@@ Acknowledgments
This work is supported by the Spanish grants PGC2022-126078NB-C21 and PID2024-160228NB-I00, funded by MCIN/AEI/10.13039/501100011033 and ‘ERDF A way of making Europe’, grant E21_23R funded by the Aragon Government and the European Union, and the NextGenerationEU Recovery and Resilience Program on ‘Astrofísica y Física de Altas Energías’ CEFCA-CAPA-ITAINNOVA. The authors would like to acknowledge the contribution of the COST Actions CA18108 “Quantum gravity phenomenology in the multi-messenger approach” and CA23130 “Bridging high and low energies in search of quantum gravity”.